OpenStudy (kendricklamar2014):

Write the expression using a single exponent. 2^8/2^7

OpenStudy (anonymous):

So in general, what happens when you divide numbers that have the same base number but have different powers, you get this,\[\frac{ x^a }{ x^b }=x^{a-b}\]can you tell what the "a" and "b" are in your problem and solve this now?

OpenStudy (kendricklamar2014):

8 and 7

OpenStudy (kendricklamar2014):

so 8-7= 1

OpenStudy (kendricklamar2014):

thanks @Embryo